OpenStudy (anonymous):

consider the line whose equation is 4x-8y=8 A.) write the equation in slope-intercept form B.) write a slope of a line perpendicular to this line C.) write the equation of the line, in point-slope form, that is perpendicular to the line that passes through the point (1, 2)

OpenStudy (triciaal):

A)slope-intercept form y =mx + b B) slope of perpendicular is -1/m (using the value of m from part A above) C) point-slope form (y-y1)=m(x-x1) the given point (!,2) is (x1,y1)